OpenStudy (anonymous):

Use the properties of logarithms to expand the following logarithms completely. log(5) ^3»(xz)

OpenStudy (anonymous):

The little arrows are a radical.

OpenStudy (psymon):

The 5 is just raised to the 3rd and not a base, correct?

OpenStudy (anonymous):

The five is the base.

OpenStudy (anonymous):

The three is a cube root of xz

OpenStudy (jdoe0001):

\(\large log_5\sqrt{3xz} \ \ \ ?\)

OpenStudy (psymon):

Gotcha. And I'm still not used to how people write out equations like jdoe just did, so that will always look funky xD

OpenStudy (jdoe0001):

\(\bf \huge log_5(\sqrt[3]{xz}) \) then

OpenStudy (anonymous):

Yes.

OpenStudy (jdoe0001):

don't forget the \( hehe

OpenStudy (anonymous):

Okay XD But anyways. Now that we have the problem written out so it makes sense, help me solve it?? XD

OpenStudy (jdoe0001):

keep in mind that \(\bf \huge a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (jdoe0001):

from there just apply the log rule for the exponents

OpenStudy (anonymous):

so it's x (z/3)????

OpenStudy (anonymous):

\[\log_{5}\sqrt[3]{xz} = \log_{5}(xz)^{1/3} = \frac{1}{3}\log_5(xz)=\frac{1}{3}(\log_5x+\log_5z)\]

OpenStudy (jdoe0001):

as shown by @walac ^

OpenStudy (jdoe0001):

the exponent comes out as coefficient, then you expand the factors inside as shown

OpenStudy (anonymous):

Ohhh, okay